There are 2 wires of same material. Their radii & lengths are both in the ratio 1: 2. If the extensions produced are equal, what is the ratio of the loads?

Grade:11

1 Answers

Vikas TU
7 years ago
let l1 be the extension in wire 1 then,
l1 = Fl/YA
and l2 be the extension in wire  2 then,
l2 = F’l’/YA’
Now Given, l1 = l2
hence,
Fl/YA = F’l’/YA’
F/F’ = Al’/A’l
       = (pi*r^2)*(2l)/(pi*4*r^2)*(l)
       = 2/4
       = ½